# Spin flip locking by the tunneling and relaxation in a driven double quantum dot with spin-orbit coupling

### v1

@@ II.3 Two-qubit representation
The first coefficient $c_{00}$ corresponds to the shift of the energy zero and does not affect the system dynamics. The coefficient $22c_{03}=(\Delta^{(1)}_{2}+\Delta^{(2)}_{2})/$ 2 is the averaged distance from (10) between the levels of the charge qubit. The coefficient $22c_{30}=(\Delta^{(1)}_{1}+\Delta^{(2)}_{1})/$ 2 is the averaged Zeeman splitting from (9) slightly corrected by the presence of SOC. The last coefficient $c_{33}$ is zero for the case of the equal Zeeman splittings in the left and right QDs and its nonzero value reflects the slight difference between the spin splitting in two dots created by the combination of SOC and detuning. This coefficient reflects the SOC-induced interaction between the spin and charge subsystems and in our modeling has a typical amplitude $|c_{33}|\sim 10^{-4}$ $\mu eV$ which allows treating SOC as a perturbation while keeping the concept of the spin and charge two-level subsystems, each described by its own relaxation parameters studied in the next Subsection. The last term $F_{4\times 4}$ in (14) is given by the decomposition of the operator $I^{(1)}_{2\times 2}\otimes f_{2}(x)$, where $I^{(1)}_{2\times 2}$ is $2\times 2$ identity matrix in the subspace of the spin qubit, and essentially depends on the shape of the driving function $f_{2}(x)$, creating in general the matrix $F_{4\times 4}$ with all non-zero elements. The notation (14) allows us to apply the standard technique of the density matrix formalism for the evolution of the system of two coupled qubits described in the next Section.
@@ II.4 Evolution of the density matrix
Since we are interested in the evolution modeling on different time scales including the ones comparable with the relaxation times, we will include the relaxation and decoherence for our system via the density matrix approach. We suppose that each of the spin and charge qubits interacts with its corresponding thermostat independently and is described by the specific phase and energy dissipators and the relaxation rates $\Gamma_{\varphi}^{(q)}$, $\Gamma_{e}^{(q)}$ for the spin and charge subsystems, providing the additive impacts into the global system dissipator Hofer2017 ; Cattaneo2019 . The full two-qubit density matrix describes the evolution of two coupled qubits governed by the equation written in the Lindblad form in the instant (adiabatic) basis Nori2023 ; Makhlin2001 ; Bastrakova2021 :
@@
We consider the approximation where each of the spin and charge subsystems interacts with its own dominating source of relaxation. For the GaAs double dots it can be the interaction with the nuclear spins for the spin subsystem and the interaction with the structural defects and phonons for the charge subsystem. Such approximation allows treating their reservoir parameters separately, introducing the phase and energy relaxation rates $\Gamma_{\varphi}^{(q)}$ and $\Gamma_{e}^{(q)}$ for each qubit subsystem ($q=1,2$) as follows Makhlin2001 :
@@
It should be mentioned that the energy relaxation rates in Eq.(II.4) are energy-dependent, i.e. they vary with the inter-level distance $\Delta\overline{E}^{(q)}$ for the adiabatic basis. In Eq.(II.4) $k_{B}$ and $\tau$ are the Boltzmann constant and the thermostat temperature, and the parameters $\eta^{(q)}$ and $\alpha^{(q)}$ describe the subsystem-thermostat coupling type and strength, respectively. For convenience in what follows we will use the characteristic relaxation times defined via (II.4) at the initial moment $t=0$: $T_{\varphi}^{(q)}=2\pi/\Gamma_{\varphi}^{(q)}(0)$ and $T_{e}^{(q)}=2\pi/\Gamma_{e}^{(q)}(0)$. The energy parameters $\Delta\overline{E}^{(q)}$ describe the inter-level distance in the instant basis,
@@
After solving Eq.(II.4) for the given parameters of the qubit subsystems and the driving in Eq.(14) we calculate the average values tracking the evolution of the each subsystem: the spin projections $\langle\sigma_{x,y,z}\rangle(t)$ tracking the state of the spin qubit where we are mainly interested in the evolution of the $z$-projection of the spin, and the electron position $\langle x/d\rangle(t)$ tracking the state of the charge qubit.
@@ III.1 EDSR affected by the spin relaxation
We consider a typical double dot formed by electrostatic gates in a GaAs nanowire similar to the ones used in the experiments Studenikin2018 and modeled in our preceding papers KS2022 ; PRB2023 . We begin with the discussion of the relaxation effects on the most basically considered pure EDSR regime. In Fig.3 we show the stroboscopic evolution of the $z$-projection of the spin $\langle\sigma_{z}\rangle(t)$ (in units of /2Planck-constant-over-$\hbar/2$) for the moments of time $t=nT$ where $T$ is the period of the driving field obtained after solving Eq.(II.4) on $5000$ $T$ time interval with the frequency matching the level distance $E_{2}-E_{1}$ (see Fig.1) defined by the Zeeman splitting $\Delta_{1}=20\mu eV$. The total evolution time is about $1$ $\mu s$. The system parameters are the same as in Subsection II.B and in Fig.2 where we choose the moderate driving amplitude $V_{d}=30$ $\mu eV$. On panel (a) the plot is shown vs time measured in microseconds (upper scale) in the driving field period (lower scale). On panel (b) the stroboscopic evolution of all three spin projections representing the spin qubit is shown on its Bloch sphere which can be introduced for both spin and charge qubit separately,
@@
Here the spin and charge qubit evolution corresponds to $q=1$ and $q=2$ in (24), respectively. The initial state is the ground spin-down state with energy $E_{1}$ in the right QD (see Fig.1), being on the south pole of the sphere in Fig.3(b). Black curve shows the coherent evolution when there is no coupling with the thermostat. Blue curve is for the relaxation times $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.4~{}\mu s$. As to the spin relaxation, we consider it to be much longer, $T_{\varphi}^{(1)}=T_{e}^{(1)}=10T_{\varphi}^{(2)}$. Red curve is for the shorter relaxation times $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.2~{}\mu s$, and again, $T_{\varphi}^{(1)}=T_{e}^{(1)}=10T_{\varphi}^{(2)}$. The thermostat temperature for the blue and red curves is $\tau=100$ $mK$. We have considered rather long relaxation times in order to illustrate the evolution on the smaller EDSR frequency in this Subsection (compared to the frequencies in the next Subsection) before they are damped completely.
Figure 3: (a) Stroboscopic evolution of $\langle\sigma_{z}\rangle(t)$ (in units of /2Planck-constant-over-$\hbar/2$) for the main EDSR harmonic at the resonance between the levels $E_{1}$ and $E_{2}$ in Fig.1. Plot is shown vs time (upper scale) and vs the number of driving field periods (lower scale); (b) Stroboscopic evolution of the vector (24) for the spin qubit ($q=1$) shown on its Bloch sphere. The system parameters are the same as in Fig.2 with the driving frequency defined by the Zeeman splitting $\Delta_{1}=20\mu eV$ and the amplitude $V_{d}=30$ $\mu eV$. The initial state is the ground spin-down state with energy $E_{1}$ in the right QD marked by the black downward arrow. Black curve is for the absence of coupling with the thermostat. Blue curve is for the relaxation times $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.4$ $\mu s$, and we take everywhere $T_{\varphi}^{(1)}=T_{e}^{(1)}=10T_{\varphi}^{(2)}$. Red curve is for $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.2$ $\mu s$. The thermostat temperature in two latter cases $\tau=100$ mK. The relaxation provides trivial damping for pure two-level spin dynamics without tunneling.
@@ III.2 Spin flip locking by the tunneling and relaxation
We continue with the evolution analysis for another type of the resonance where the spin flip is accompanied by the effective tunneling. This type of resonance corresponds to the pair of levels $E_{1}$ and $E_{4}$ in Fig.1 where the spin flip is accompanied with the effective interdot tunnelingKS2022 ; PRB2023 . Here the basic system parameters are the same as in the previous Subsection. The frequency of the resonance between the levels $E_{1}$ and $E_{4}$ is determined by the sum $\Delta_{1}+\Delta_{2}=75$ $\mu eV$ giving us the linear frequency $f\sim 18~{}GHz$ which is still within the frequency bands used in the experiments Studenikin2018 . The characteristic timescale for the developing spin and coordinate dynamics is much faster than in the previous Subsection, it is in between of the charge and spin relaxation timescales which allows to capture well-established spin rotations. One should mention that the amplitude of the driving field $V_{d}=30\mu eV$ is lower than the detuning amplitude, meaning that we are not in the effective LZSM regime with the level anticrossing Nori2023 ; KS2022 but purely in the resonance at $E_{4}-E_{1}=\omega$. The initial state for the evolution is again the ground state $E_{1}$. The total evolution time for the modeling is about 100 $ns$ which is within the range of the earlier experiments on EDSR-induced current oscillations in GaAs gate-defined double dots where no significant damping has been observed Nowack2007 .
@@
Figure 4: Evolution of (a) the mean spin projection $\langle\sigma_{z}\rangle(t)$ and (b) the mean coordinate $\langle x\rangle(t)$ (in units of $d$) for the spin-flip tunneling resonance between the levels $E_{1}$ and $E_{4}$ on the frequency close to the combined splitting $\Delta_{1}+\Delta_{2}=75$ $\mu eV$ and accompanied by the quick relaxation between the levels $E_{4}$ and $E_{2}$ in Fig.1. The evolution is plotted vs time (upper scale) and the number of driving field periods (lower scale). All the other parameters are the same as in Fig.3. The stabilization of the flipped spin in the right QD can be observed due to the fast charge relaxation and slow spin relaxation.
Figure 5: Evolution of the vector (24) on the Bloch sphere for (a) the spin qubit and (b) the charge qubit. The initial state is on the south pole for both qubits. The type of the resonance and the parameters correspond to Fig.4. The final position of the vector (24) is on the north pole for the spin qubit in panel (a) and on the south pole for the charge qubit in panel (b).
@@
In Fig.4 we show the evolution for the mean values of the spin projection $\langle\sigma_{z}\rangle(t)$ and the coordinate $\langle x\rangle(t)$, the latter measured in units of $d$, obtained after solving Eq.(II.4) on the $2000$ $T$ with the frequency matching the level distance $E_{4}-E_{1}$ (see Fig.1). The associated evolution on the Bloch sphere of the vector (24) is shown in Fig.5 for (a) spin qubit and (b) charge qubit. The initial state is on the south pole for both qubits. It is clear that for all considered values of the relaxation parameters the evolution on the Bloch sphere is similar for both qubits and can be described as a small-radius rotation around the $z$ axis with the full-scale oscillations between the south and the north poles. The final position of the vector (24) in Fig.5, as it can be concluded from Fig.4, is on the north pole for the spin qubit in panel (a) and on the south pole for the charge qubit in panel (b).
@@
As in the previous Subsection, we consider three sets of the relaxation parameters. For the first set corresponding to the black curves there is no coupling with the thermostat, i.e. it is the basic case of the coherent evolution studied earlier KS2022 ; PRB2023 . For the second set corresponding to the blue curves the charge qubit relaxation times $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.4$ $\mu s$, and the spin qubit relaxation times are ten times slower, $T_{\varphi}^{(1)}=T_{e}^{(1)}=10T_{\varphi}^{(2)}$. This is a typical situation in semiconductor nanowires and quantum dots where the spin relaxation is substantially slower than the momentum relaxation. The spin relaxation time in the GaAs-based double dot system with the detuning is expected to be much longer, at least an order of magnitude Vandersypen2013 ; Vandersypen2014 . For the third set described by the red curves the relaxation is faster, $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.2$ $\mu s$, and $T_{\varphi}^{(1)}=T_{e}^{(1)}=10T_{\varphi}^{(2)}$. The fastest charge relaxation time of $200$ $ns$ is of the same order as the maximum evolution time of $700$ $ns$ in Fig.4 and is comparable to the coherence time obtained in recent calculations for the double dots with detuning Krzywda2021 . The considered scale of spin and charge relaxation times means that the charge (coordinate) subsystem evolves during the evolution towards its ground state which is the lowest level $E_{1}$ in the right QD while the spin subsystem is still far from its ground state. The thermostat temperature for the second and third set is set at $\tau=100$ $mK$ being in agreement with the typical experimental conditions Studenikin2018 . Such temperature corresponds to the thermal energy of about $8.75\mu eV$ being lower that all the other energy parameters of the model, meaning that the thermal-induced jumps to the higher levels have low probability.
@@
The most striking feature in Fig.4 is the observed spin flip stabilization, or spin flip locking in the right QD when the relaxation is introduced in combination with the interdot tunneling. This effect can be explained by comparing the fast charge relaxation with the slow spin relaxation and the resonant driving. Namely, the driving on the resonance $E_{4}-E_{1}=\omega$ excites the spin-flip transitions accompanied with the interdot tunneling from right to the left QD. This is illustrated in Fig.6 where an example of the level occupancies $|C_{n}(t)|^{2}$ for the levels $E_{1},\ldots,E_{4}$ from Fig.1 is shown for the case of strong relaxation labeled by the red curve in Fig.4. The main purpose of Fig.6 is to show the evolution in the basis of the coordinate- and spin-resolved states of the static Hamiltonian (2) which is readily connected to the evolution of the mean values in Fig.4. In Fig.6 one can see that for the resonance between the levels $E_{1}$ and $E_{4}$ the off-resonance spin-up level $E_{2}$ occupation $|C_{2}(t)|^{2}$ monotonically grows with time until a certain saturation of the spin flip in the right QD is achieved. The spin-down level $E_{3}$ population $|C_{3}(t)|^{2}$ in the left QD demonstrates the slow-amplitude steady behavior while both of remaining level populations $|C_{1}(t)|^{2}$ in the left QD and $|C_{4}(t)|^{2}$ in the right QD oscillate in resonance and demonstrate a decaying behavior. It is evident that a steady state in terms of the level populations is formed on longer times with the stable spin flip in the right QD. We would like to stress that the considered example of the dynamics in Fig.4 - Fig.6 demonstrates how useful the second (charge) qubit can be for the creation of the locked population of the excited state of the first (spin) qubit, finally leaving the second qubit in its ground state in the right QD.
@@
Figure 7: The same as in Fig.4 for the second subharmonic $s=2$ in (25) of the $E_{1}$ $E_{4}$ resonant transition in Fig.1. The type of the resonance and the parameters (except the lower driving frequency and longer evolution time) are the same as in Fig.4. The relaxation leads to the dampening of the spin and coordinate oscillations and to the stabilization of the flipped spin.

### v2

@@ II.3 Two-subsystem representation
The first coefficient $c_{00}$ corresponds to the shift of the energy zero and does not affect the system dynamics. The coefficient $22c_{03}=(\Delta^{(1)}_{2}+\Delta^{(2)}_{2})/$ 2 is the averaged distance from (10) between the levels of the charge subsystem. The coefficient $22c_{30}=(\Delta^{(1)}_{1}+\Delta^{(2)}_{1})/$ 2 is the averaged Zeeman splitting from (9) slightly corrected by the presence of SOC. The last coefficient $c_{33}$ is zero for the case of the equal Zeeman splittings in the left and right QDs and its nonzero value reflects the slight difference between the spin splitting in two dots created by the combination of SOC and detuning. This coefficient reflects the SOC-induced interaction between the spin and charge subsystems and in our modeling has a typical amplitude $|c_{33}|\sim 10^{-4}$ $\mu eV$ which allows treating SOC as a perturbation while keeping the concept of the spin and charge two-level subsystems, each described by its own relaxation parameters studied in the next Subsection. The last term $F_{4\times 4}$ in (14) is given by the decomposition of the operator $I^{(1)}_{2\times 2}\otimes f_{2}(x)$, where $I^{(1)}_{2\times 2}$ is $2\times 2$ identity matrix in the subspace of the spin subsystem, and essentially depends on the shape of the driving function $f_{2}(x)$, creating in general the matrix $F_{4\times 4}$ with all non-zero elements. By analyzing the representation (14) one can see that the spin subsystem is described by the third term with the level splitting given by $c_{30}$ while the charge subsystem is described by the second term in (14) with the level splitting given by the $c_{03}$. Both the spin and charge subsystems in the representation (14) do not belong to a single dot but to the ground four-level manifold of the whole double dot structure. After being introduced, the notation (14) allows us to apply the standard technique of the density matrix formalism for the evolution of the system of two coupled subsystems described in the next Section.
@@ II.4 Evolution of the density matrix
Since we are interested in the evolution modeling on different time scales including the ones comparable with the relaxation times, we will include the relaxation and decoherence for our system via the density matrix approach. We suppose that each of the spin and charge subsystems interacts with its corresponding thermostat independently and is described by the specific phase and energy dissipators and the relaxation rates $\Gamma_{\varphi}^{(q)}$, $\Gamma_{e}^{(q)}$ for the spin and charge subsystems, providing the additive impacts into the global system dissipator Hofer2017 ; Cattaneo2019 . The full two-subsystem density matrix describes the evolution of two coupled subsystems governed by the equation written in the Lindblad form in the instant (adiabatic) basis Nori2023 ; Makhlin2001 ; Bastrakova2021 :
@@
We consider the approximation where each of the spin and charge subsystems interacts with its own dominating source of relaxation. For the GaAs double dots it can be the interaction with the nuclear spins for the spin subsystem and the interaction with the structural defects and phonons for the charge subsystem. This approximation allows treating their reservoir parameters separately, introducing the phase and energy relaxation rates $\Gamma_{\varphi}^{(q)}$ and $\Gamma_{e}^{(q)}$ for each subsystem ($q=1,2$) as follows Makhlin2001 :
@@
It should be mentioned that the energy relaxation rates in Eq.(II.4) are energy-dependent, i.e. they vary with the inter-level distance $\Delta\overline{E}^{(q)}$ for the adiabatic basis. In Eq.(II.4) $k_{B}$ and $\tau$ are the Boltzmann constant and the thermostat temperature, and the parameters $\eta^{(q)}$ and $\alpha^{(q)}$ describe the subsystem-thermostat coupling type and strength, respectively. Despite the time dependence of the rates (II.4), we need a parameter reference for the relaxation times for the spin and coordinate subsystem. To introduce this, we will use the characteristic relaxation times defined via (II.4) at the initial moment $t=0$: $T_{\varphi}^{(q)}=2\pi/\Gamma_{\varphi}^{(q)}(0)$ and $T_{e}^{(q)}=2\pi/\Gamma_{e}^{(q)}(0)$. The energy parameters $\Delta\overline{E}^{(q)}$ describe the inter-level distance in the instant basis,
@@
After solving Eq.(II.4) for the given parameters of the subsystems and the driving in Eq.(14) we calculate the average values tracking the evolution of each subsystem: the spin projections $\langle\sigma_{x,y,z}\rangle(t)$ tracking the state of the spin subsystem where we are mainly interested in the evolution of the $z$-projection of the spin, and the electron position $\langle x/d\rangle(t)$ tracking the state of the charge subsystem.
@@ III.1 EDSR affected by the spin relaxation
We consider a typical double dot formed by electrostatic gates in a GaAs nanowire similar to those used in the experiments Studenikin2018 and modeled in our preceding papers KS2022 ; PRB2023 . We begin with a discussion of the relaxation effects on the most basically considered pure EDSR regime. In Fig.3 we show the stroboscopic evolution of the $z$-projection of the spin $\langle\sigma_{z}\rangle(t)$ (in units of /2Planck-constant-over-$\hbar/2$) for the moments of time $t=nT$ where $T$ is the period of the driving field obtained after solving Eq.(II.4) on $5000$ $T$ time interval with the frequency matching the level distance $E_{2}-E_{1}$ (see Fig.1) defined by the Zeeman splitting $\Delta_{1}=20\mu eV$. The total evolution time is about $1$ $\mu s$. The system parameters are the same as in Subsection II.B and in Fig.2 where we choose the moderate driving amplitude $V_{d}=30$ $\mu eV$. On panel (a) the plot is shown vs time measured in microseconds (upper scale) in the driving field period (lower scale). On panel (b) the stroboscopic evolution of all three spin projections representing the spin subsystem is shown on its Bloch sphere which can be introduced for both spin and charge subsystem separately,
@@
Here the spin and charge subsystem evolution corresponds to $q=1$ and $q=2$ in (24), respectively. The initial state is the ground spin-down state with energy $E_{1}$ in the right QD (see Fig.1), being on the south pole of the sphere in Fig.3(b). Black curve shows the coherent evolution when there is no coupling with the thermostat. Blue curve is for the relaxation times $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.4~{}\mu s$. As to the spin relaxation, we consider it to be much longer, $T_{\varphi}^{(1)}=T_{e}^{(1)}=10T_{\varphi}^{(2)}$. Red curve is for the shorter relaxation times $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.2~{}\mu s$, and again, $T_{\varphi}^{(1)}=T_{e}^{(1)}=10T_{\varphi}^{(2)}$. The thermostat temperature for the blue and red curves is $\tau=100$ $mK$. We have considered rather long relaxation times in order to illustrate the evolution on the smaller EDSR frequency in this Subsection (compared to the frequencies in the next Subsection) before they are damped completely.
Figure 3: (a) Stroboscopic evolution of $\langle\sigma_{z}\rangle(t)$ (in units of /2Planck-constant-over-$\hbar/2$) for the main EDSR harmonic at the resonance between the levels $E_{1}$ and $E_{2}$ in Fig.1. Plot is shown vs time (upper scale) and vs the number of driving field periods (lower scale); (b) Stroboscopic evolution of the vector (24) for the spin subsystem ($q=1$) shown on its Bloch sphere. The system parameters are the same as in Fig.2 with the driving frequency defined by the Zeeman splitting $\Delta_{1}=20\mu eV$ and the amplitude $V_{d}=30$ $\mu eV$. The initial state is the ground spin-down state with energy $E_{1}$ in the right QD marked by the black downward arrow. Black curve is for the absence of coupling with the thermostat. Blue curve is for the relaxation times $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.4$ $\mu s$, and we take everywhere $T_{\varphi}^{(1)}=T_{e}^{(1)}=10T_{\varphi}^{(2)}$. Red curve is for $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.2~{}\mu s$. The thermostat temperature in two latter cases $\tau=100$ mK. The relaxation provides trivial damping for pure two-level spin dynamics without tunneling.
@@ III.2 Spin flip locking by the tunneling and relaxation
We continue with the evolution analysis for another type of resonance for the pair of levels $E_{1}$ and $E_{4}$ in Fig.1 where the spin flip is accompanied by the effective interdot tunnelingKS2022 ; PRB2023 . Here the basic system parameters are the same as in the previous Subsection. The frequency of the resonance between the levels $E_{1}$ and $E_{4}$ is determined by the sum $\Delta_{1}+\Delta_{2}=75$ $\mu eV$ giving us the linear frequency $f\sim 18~{}GHz$. It is still within the frequency bands used in the experiments Studenikin2018 . The characteristic timescale for the developing spin and coordinate dynamics is much faster than in the previous Subsection. It is in between of the charge and spin relaxation timescales which allows to capture well-established spin rotations. One should mention that the amplitude of the driving field $V_{d}=30\mu eV$ is lower than the detuning amplitude, meaning that we are not in the effective LZSM regime with the level anticrossing Nori2023 ; KS2022 but purely in the resonance at $E_{4}-E_{1}=\omega$. The initial state for the evolution is again the ground state $E_{1}$. The total evolution time for the modeling is about 100 $ns$ which is within the range of the earlier experiments on EDSR-induced current oscillations in GaAs gate-defined double dots where no significant damping has been observed Nowack2007 .
@@
Figure 4: Evolution of (a) the mean spin projection $\langle\sigma_{z}\rangle(t)$ and (b) the mean coordinate $\langle x\rangle(t)$ (in units of $d$) for the spin-flip tunneling resonance between the levels $E_{1}$ and $E_{4}$ on the frequency close to the combined splitting $\Delta_{1}+\Delta_{2}=75$ $\mu eV$ and accompanied by the quick relaxation between the levels $E_{4}$ and $E_{2}$ in Fig.1. The evolution is plotted vs time (upper scale) and the number of driving field periods (lower scale). All the other parameters and the labeling of the curves are the same as in Fig.3. The stabilization of the flipped spin in the right QD can be observed due to the fast charge relaxation and slow spin relaxation. The insets show the evolution of the vector (24) on the Bloch sphere for (a) the spin subsystem and (b) the charge subsystem. The initial state is on the south pole for both subsystems.
@@
In Fig.4 we show the evolution for the mean values of the spin projection $\langle\sigma_{z}\rangle(t)$ and the coordinate $\langle x\rangle(t)$, the latter measured in units of $d$, obtained after solving Eq.(II.4) on the $2000$ $T$ with the frequency matching the level distance $E_{4}-E_{1}$ (see Fig.1). The associated evolution on the Bloch sphere of the vector (24) is shown in the insets on Fig.4 for (a) spin subsystem and (b) charge subsystem. The initial state is on the south pole for both subsystems. It is clear that for all considered values of the relaxation parameters the evolution on the Bloch sphere is similar for both subsystems and can be described as a small-radius rotation around the $z$ axis with the full-scale oscillations between the south and the north poles. The final position of the vector (24) in the insets of Fig.4, as it can be concluded from the plots in Fig.4, is on the north pole for the spin subsystem in panel (a) and on the south pole for the charge subsystem in panel (b).
@@
As in the previous Subsection, we consider three sets of the relaxation parameters. For the first set corresponding to the black curves there is no coupling with the thermostat, i.e. it is the basic case of the coherent evolution studied earlier KS2022 ; PRB2023 . For the second set corresponding to the blue curves the charge subsystem relaxation times $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.4$ $\mu s$, and the spin subsystem relaxation times are ten times slower, $T_{\varphi}^{(1)}=T_{e}^{(1)}=10T_{\varphi}^{(2)}$. This is a typical situation in semiconductor nanowires and quantum dots where the spin relaxation is substantially slower than the momentum relaxation. The spin relaxation time in the GaAs-based double dot system with the detuning is expected to be much longer, at least an order of magnitude Vandersypen2013 ; Vandersypen2014 . For the third set described by the red curves the relaxation is faster, $T_{\varphi}^{(2)}=T_{e}^{(2)}=0.2$ $\mu s$, and $T_{\varphi}^{(1)}=T_{e}^{(1)}=10T_{\varphi}^{(2)}$. The fastest charge relaxation time of $200$ $ns$ is of the same order as the maximum evolution time of $700$ $ns$ in Fig.4 and is comparable to the coherence time obtained in recent calculations for the double dots with detuning Krzywda2021 . The considered scale of spin and charge relaxation times means that the charge (coordinate) subsystem evolves during the evolution towards its ground state which is the lowest level $E_{1}$ in the right QD while the spin subsystem is still far from its ground state. The thermostat temperature for the second and third cases is set at $\tau=100$ $mK$ being in agreement with the typical experimental conditions Studenikin2018 . Such temperature corresponds to the thermal energy of about $8.75\mu eV$ being lower than all the other energy parameters of the model, meaning that the thermal-induced jumps to the higher levels have low probability.
@@
The most striking feature in Fig.4 is the observed spin flip stabilization, or spin flip locking in the right QD when the relaxation is introduced in combination with the interdot tunneling. This effect can be explained by comparing the fast charge relaxation with the slow spin relaxation and the resonant driving. Namely, the driving on the resonance $E_{4}-E_{1}=\omega$ excites the spin-flip transitions accompanied by the interdot tunneling from the right to the left QD. This is illustrated in Fig.5 where an example of the level occupancies $|C_{n}(t)|^{2}$ for the levels $E_{1},\ldots,E_{4}$ from Fig.1 is shown for the case of strong relaxation labeled by the red curve in Fig.4. The main purpose of Fig.5 is to show the evolution in the basis of the coordinate- and spin-resolved states of the static Hamiltonian (2) which is readily connected to the evolution of the mean values in Fig.4. In Fig.5 one can see that for the resonance between the levels $E_{1}$ and $E_{4}$ the off-resonance spin-up level $E_{2}$ occupation $|C_{2}(t)|^{2}$ monotonically grows with time until a certain saturation of the spin flip in the right QD is achieved. The populations $|C_{1}(t)|^{2}$ and $|C_{4}(t)|^{2}$ oscillate with the associated Rabi frequency for this two-level resonance modulated by the decaying envelope function while the spin-down level $E_{3}$ population $|C_{3}(t)|^{2}$ in the left QD demonstrates the slow-amplitude steady behavior. The observed Rabi frequency is higher for the $E_{1}$ $E_{4}$ resonance than for the EDSR resonance considered in the previous Subsection which has been observed in our previous papers KS2022 ; PRB2023 since the former involves the greater spatial interdot displacement leading to a higher matrix element amplitude for the SOC. As to the stable and slow growth of $|C_{2}(t)|^{2}$, it is a result of the two processes: the fast interdot spin flip tunneling and the slow interdot charge relaxation. As a result, from Fig.5 it is evident that a steady state in terms of the level populations is formed on longer times with the stable spin flip in the right QD. We would like to stress that the considered example of the dynamics in Fig.4 - Fig.5 demonstrates how useful the second (charge) subsystem can be for the creation of the locked population of the excited state of the first (spin) subsystem, finally leaving the second subsystem in its ground state in the right QD.
@@
Figure 6: The same as in Fig.4 for the second subharmonic $s=2$ in (25) of the $E_{1}$ $E_{4}$ resonant transition in Fig.1. The labeling of the curves is the same as in Fig.3 and Fig.4. The type of the resonance and the parameters (except the lower driving frequency and longer evolution time) are the same as in Fig.4. The relaxation leads to the dampening of the spin and coordinate oscillations and to the stabilization of the flipped spin.